OpenStudy (anonymous):

Solve cos x tan x=1/2 for x over the interval [0,2 pi). a. pi/6 b. pi/3 and 5pi/3 c. pi/6 and 11pi/6 d. pi/6 and 5pi/6 please help me:(

OpenStudy (anonymous):

\[\cos x \tan x=\frac{ 1 }{ 2 },\cos x \frac{ \sin x }{ \cos x }=\frac{ 1 }{ 2 }\] \[\sin x=\frac{ 1 }{ 2 }=\sin \frac{ \pi }{ 6 },\sin \left( \pi-\frac{ \pi }{ 6 } \right),x=\frac{ \pi }{ 6 },\frac{ 5 \pi }{ 6 }\]

OpenStudy (bradely):

cos x tan x=1/2 tan x can be written as sin x /cos x. cos x (tan x)=1/2 cos x (sin x/cos x) =1/2 sin x =1/2 x = pi/6 General solution: x = npi+(-1)^n(pi/6) At n = 0 x =pi/6 At n = 1 x =pi -(pi/6) =5pi/6 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

my teacher tried explaining that to me and i could never understand it but thanks i have a pretty ok idea about it...THANKS!!!!